OpenStudy (anonymous):

Can somebody walk this through with me? If the side length of a square can be represented by 2x + 10 and its area is 256 square units, find the value of x.

OpenStudy (mathstudent55):

The area of a square is the square of the side, right?

OpenStudy (anonymous):

Yeah

OpenStudy (mathstudent55):

This square has a side of 2x + 10, so its area is \((2x + 10)^2 \)

OpenStudy (anonymous):

2x = 256- 10 2x = 246 x = 123 that easy to do!!

OpenStudy (mathstudent55):

We also know the area is 256 sq units.

OpenStudy (anonymous):

It's not 123, that's what I wrote and it was marked incorrect.

OpenStudy (mathstudent55):

Not quite, @ashuu

OpenStudy (mathstudent55):

Since we have an expression for the area and the area itself, we make them equal and solve the equation.

OpenStudy (anonymous):

that nt r8 !!

OpenStudy (mathstudent55):

\( (2x + 10 )^2 = 256 \)

OpenStudy (anonymous):

@ashuu you saved like four letters typing that

OpenStudy (anonymous):

okey i havt seen that mathstudent55

OpenStudy (mathstudent55):

Square the binomial on the left side: \(4x^2 + 40x + 100 = 256 \)

OpenStudy (mathstudent55):

Now subtract 256 from both sides. \(4x^2 + 40x - 156 = 0 \)

OpenStudy (anonymous):

4x2+40x=156 x2+x=39 The two means squared

OpenStudy (mathstudent55):

We need to factor the left side, but we can divide both sides by 4 since every term is divisible by 4. \(x^2 + 10x - 39 = 0 \)

OpenStudy (anonymous):

Yeah so do you do the square root of 39

OpenStudy (anonymous):

Correct?

OpenStudy (anonymous):

The final answer is supposed to be x=3. I am still really confused as to how you get to that.

OpenStudy (mathstudent55):

I'm getting there. Just follow along. Now we can factor the left side: \( (x - 3)(x + 13) = 0 \) Set each factor equal to zero: \( x - 3 = 0\) or \(x + 13 = 0 \) Solve each equation: \( x = 3\) or \(x = -13 \)

OpenStudy (anonymous):

They both end up equaling zero. Whats the next step?

OpenStudy (anonymous):

What are you writing that seriously takes 4 minutes to type?

OpenStudy (mathstudent55):

Now that we have two solutions, x = 3 and x = -13. We try them in our original problem. The side of the square is 2x + 10: For x = 3, 2x + 10 = 2(3) + 10 = 6 + 10 = 16 For x = -13, 2x + 10 = 2(-13) + 10 = -26 + 10 = -16 Now we see that using x = 3, we get a side length of 16 units. A positive number is certainly acceptable as a side length. We keep the answer x = 3. We also see that using the answer x = -13 gives us a side length of -16 units. A side length cannot be a negative number, so we must disregard the x = -13 solution. The answer to the problem is: x = 3

OpenStudy (anonymous):

Okay thanks

OpenStudy (mathstudent55):

Let me quote you: "Can somebody walk this through with me?" If you want a walk through, you need to be patient. I could have done all the math with no explanations much faster, but then I'd understand it, but you wouldn't. That wouldn't be very helpful.

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

Yeah sorry I got impatient, my Dad is bugging me to go to bed.

OpenStudy (mathstudent55):

listen to your dad. do some more math tomorrow